OpenStudy (anonymous):

HEY, im having trouble with converting x to the power 3 divided by 3 into a single um. for example, I know that 3 over x to the power three would be 3x to the power minus three, but how would you write x to the power three over three?

OpenStudy (skullpatrol):

Could you write that in symbols please?

OpenStudy (anonymous):

\[^{x^3}/3 = ? \]

OpenStudy (anonymous):

I know that \[3/x^3 = 3x^-3 \]

OpenStudy (anonymous):

im sorry I don't know how to use this equation thingy

OpenStudy (skullpatrol):

$$\huge \frac{3}{x^3}=3x^{-3}$$

OpenStudy (anonymous):

I know:P im asking what \[x^3/3 \] is

OpenStudy (skullpatrol):

$$\huge \frac{x^3}{3}=\frac{1}3x^{3}$$

OpenStudy (skullpatrol):

$$\huge \frac{a}{b}=a*\frac{1}{b}$$

OpenStudy (anonymous):

ooh, thankyou:)

OpenStudy (skullpatrol):

$$\huge \frac{1}{3}=3^{-1}$$